# Flexible cost-penalized Bayesian model selection: developing inclusion paths with an application to diagnosis of heart disease

### v2

@@ 1 Introduction
When applying the FND prior to other data sets, we have found that the FND prior can lead to selection of less costly models when predictor costs differ, but the cost penalization is not appropriate or sufficient for all sample sizes. In fact, at large sample sizes, the cost penalization imposed by the FND prior greatly diminishes, often causing the resulting cost-penalized model selection to closely resemble that of a standard benefit-only analysis. To establish this phenomenon, we calculated the Kullback-Leibler (KL) divergence, which measures the difference between two probability distributions, between the sets of posterior model probabilities for all candidate models produced by cost-penalized analysis with the FND prior and the benefit-only selection approach as sample size increases for several simulated data sets. Since we consider the case of linear logistic regression, for which there is no closed form integrated likelihood, we examine the KL divergence comparing posterior model probabilities from the two approaches empirically. We generated $10$ data sets of size $n=150$ with $p=9$ total candidate predictors with different costs as detailed in Section 3.1 and calculated the KL divergence between the two sets of posterior model probabilities produced by the benefit-only and FND selection approaches for each data set. Then, to imitate collection of additional data, we recursively added $300$ observations to the initial $10$ data sets, obtaining model selection results and calculating KL divergence between the posterior model probabilities produced by the two selection approaches for each new (larger) data set. Figure 2 plots the KL divergence between the posterior model probabilities produced by the two approaches as sample size increases for the $10$ collections of data sets increasing in size.
@@ 2.2 Cost-penalizing model selection
Following Fouskakis et al. (2009a), we consider the linear logistic regression setting in this paper, later applying it to the binary diagnosis of heart disease. For convenience, we use the notation of Fouskakis et al. (2009a) for indicator functions and predictor costs. We use a linear logistic regression model where $Y_{i}\in\{0,1\}$ and $X_{ij}$ denotes the $j{\color[rgb]{0,0,0}{}^{th}}$ predictor for observation $i$, where $i=1,\ldots,n$ and $j=0,\ldots,p$. Let $\gamma_{j}$ be an indicator that is equal to $1$ if predictor $X_{j}$ is included in the model and 00 if it is not. Then $\boldsymbol{\gamma}$ is a length $p$ vector of 00’s and $1$’s indicating whether each of the $p$ predictors are in the model or not. Note that the intercept is included in every model so $X_{i0}=1$ and $\gamma_{0}=1$ for all models. Then the Bayesian modeling framework is
@@
Equation (6) uses the marginal cost $\{c_{j},j=1,\ldots p\}$ of each candidate predictor, which is the cost of collecting that predictor for a single observation. When adhering to an overall budget, the practitioner may also want to consider the total cost per observation or the total cost of the model. The total cost per observation is equal to the sum of the marginal cost of all predictors in the model, i.e. $\sum_{j=1}^{p}{\color[rgb]{0,0,0}c_{j}}\gamma_{j}$, and the total cost of the model for the given observations is $n\times\sum_{j=1}^{p}{\color[rgb]{0,0,0}c_{j}}\gamma_{j}$.
@@ 2.3 Adjusted cost-penalizing functions
We propose functions $g(c_{j}/c_{0},{\color[rgb]{0,0,0}b})$ to adjust the cost ratio according to tuning parameter $b$ that satisfy the following properties:
@@
    $b=0$ implies $g(c_{j}/c_{0},{\color[rgb]{0,0,0}b})=1$ for all $j=1,\ldots,p$, reducing the prior to the uniform model prior (7) and resulting in a benefit-only model selection.
@@
    $b=1$ makes $g(c_{j}/c_{0},{\color[rgb]{0,0,0}b})$ equal to $c_{j}/c_{0}$ and reproduces the FND prior as seen in Equation (6).
@@
    When $b>1$, $g(c_{j}/c_{0},{\color[rgb]{0,0,0}b})$ penalizes predictors with costs $c_{j}>c_{0}$ more highly than the FND prior in (6). When $10<b<$ 1, $g(c_{j}/c_{0},{\color[rgb]{0,0,0}b})$ penalizes predictors with costs $c_{j}>c_{0}$ less than the FND prior but more than a benefit-only analysis. Higher values of $b$ increase $g(c_{j}/c_{0},{\color[rgb]{0,0,0}b})$ and the resulting cost penalization, leading to lower cost-adjusted prior inclusion probabilities for predictors with costs above the baseline.
@@
    When $c_{j}=c_{0}$, the $j{\color[rgb]{0,0,0}{}^{th}}$ candidate predictor cost is the same as the baseline cost. Changing $b$ does not introduce/increase penalization on any of the predictors with baseline cost.
@@ 2.4 Inclusion paths
Using the ECP and varying $b$, we create an inclusion path for each candidate predictor. To do this, we compute and plot cost-adjusted posterior inclusion probabilities for each of the candidate predictors across multiple values of $b$. See the Supplementary Material for detailed steps for constructing the inclusion path. The inclusion path visualizes a probabilistic way to learn the order in which the candidate predictors would be chosen or discarded, e.g. according to some threshold, if a different magnitude of cost penalization is used. The costs $\{c_{j},j=1,\ldots p\}$ remain unchanged; $g(c_{j}/c_{0},{\color[rgb]{0,0,0}b})$ changes how heavily candidate predictors with larger costs are penalized a priori. Our inclusion path technique is inspired by the interpretable path diagrams like LASSO and ridge (Tibshirani, 1996; Hoerl and Kennard, 1970). From the plot of inclusion paths, the practitioner can all at once study candidate predictors’ posterior inclusion probabilities for the benefit-only analysis and cost-adjusted posterior inclusion probabilities for the FND prior and a wide range of the ECP with differing $b$ values. This is because we formulated the ECP to satisfy properties (a)-(d) so that the established uniform priors and the FND prior can be easily recreated from the ECP.
@@ 3.3 Inclusion paths using adjusted cost penalization
From Figure 2(a), the posterior inclusion probabilities for $8$ of the $9$ predictors decrease as the cost penalization is increased via larger values of $b$. The predictor with the larger effect size and baseline cost maintains a posterior inclusion probability (benefit-only or cost-adjusted) at $1$ across all values of $b$, as it is not penalized a priori by the FND or ECP priors, according to property (d) from Section 2.3. The posterior inclusion probabilities for the cheap and expensive null predictors are $0.09$ and $0.25$, respectively, in the benefit-only analysis corresponding to $b=0$. The values of $10<b<$ 1 introduce cost penalization that is less severe than that from the FND prior but higher than the (nonexistent) cost penalization from the benefit-only analysis. For example, the cost-adjusted posterior inclusion probabilities for the cheap and expensive null predictors drop to $0.04$ and $0.06$ when $b=0.2$ is used with the ECP. Then the null predictor with expensive cost has 00 cost-adjusted posterior inclusion probability for $b\geq 0.7$. Thus, even slightly penalizing these null predictors for having costs above the baseline helps to move their inclusion probabilities to 00. For $b=1$ (the FND approach), the null predictor with cheap cost has cost-adjusted posterior inclusion probability $0.0002$, which decreases to 00 for $b\geq 1.1$. The posterior inclusion probability of the null predictor with baseline cost is $0.12$ in the benefit-only analysis and the cost-adjusted posterior inclusion probability equals $0.05$ for $b\geq 3.5$, as this predictor has a null effect but is not penalized by any variation of the ECP. When determining which predictors are selected with each variation of the ECP, we can consider comparing the cost-adjusted posterior inclusion probabilities for the remaining non-null predictors to a threshold, e.g. $0.5$ as in the median probability model (Barbieri and Berger, 2004). The inclusion path allows us to see which predictors’ cost-adjusted inclusion probabilities meet the desired threshold for the different versions of cost-penalized selection at one time. For example, in the benefit-only analysis, which corresponds to the ECP with $b=0$, every one of the $6$ non-null predictors have posterior inclusion probabilities beyond the $0.5$ threshold under consideration. Using the FND prior, which corresponds to the ECP with $b=1$, the cost-adjusted posterior inclusion probabilities for all non-null predictors except for the expensive one with smaller effect size exceed the $0.5$ threshold. Figure 3 indicates that the cost-adjusted posterior inclusion probabilities fall below 0.5 at b = 1, 1.1, 1.5, and 1.8 for these predictors in the following order: expensive with smaller effect size, expensive with larger effect size, cheap with smaller effect size, and cheap with larger effect size.
@@
Figure 2(b) plots the cost of the selected model for each value of $b$ used to create the inclusion path in Figure 2(a). For this data set, the MAP model and median probability model are the same for all values of $b$. We can see that the cost of the selected model drops each time a predictor is no longer in the model, e.g. inclusion falls below 0.5. Figure 2(c) plots the accuracy in the form of the C-statistic of the selected model for each value of $b$. Similar to the cost in Figure 2(b), the C-statistic drops each time a predictor leaves the selected model as a result of a higher cost penalty. A practitioner, patient, or other stakeholder may use Figures 2(b) and 2(c) in conjunction with the inclusion path so they may keep any necessary budget, or desired efficiency or accuracy in mind when selecting cost-effective predictors.
@@ 3.4 Case study: selecting cost-effective predictors to model diagnosis of heart disease
We performed Bayesian model selection using the heart disease data, applying a benefit-only approach and cost-penalized model selection as described in Section 2 with the FND and ECP priors assigned to the model space. Benefit-only model selection with uniform priors (7) on the model space selects the model with the following predictors: sex, chest pain, resting BP, ST depression, peak ST segment, major vessels, and defect type with corresponding posterior model probability $0.072$, up from a prior model probability equal to 0.000122 and with a total cost of \$$\$381.40$ per patient. Model selection using the FND prior on the model space selects the model with the four baseline predictors (sex, age, chest pain, and resting BP), with corresponding cost-adjusted posterior model probability equal to 0.5. This model that contains only baseline predictors has prior model probability equal to $0.063$ using the FND prior, and a cost equal to \$$\$4.00$ per patient. While cost-penalized selection with the FND prior leads to a less costly model that is only $2.3\%$ of the cost of the benefit-only model, this model performs worse than more costly models in terms of classification. Figure 5 displays ROC curves and corresponding costs for models selected using different values of $b$ with ECP; the values of $b$ listed correspond to the median probability model. Detection of a health condition such as heart disease is critical for doctors to be able to provide effective treatment and advice to affected patients. The cost \$$\$4.00$ model might not have acceptable false positive rates for the hospital to trust the diagnosis, and thus it is desirable to be able to adjust the cost penalization, particularly between $10<b<$ 1, for these data. Our method is key for striking a balance between the benefit-only approach and the FND approach, by furnishing a spectrum of cost penalization choices. The C-statistics for the models considered in Figure 5 are 0.934, 0.931, 0.917, 0.887, and 0.849 for the median probability models from the ECP with $b=$ 0, 0.15, 0.25, 0.35, and 1, respectively. To assess out-of-sample predictive accuracy, we also calculated a leave-one-out C-statistic for each model by obtaining predicted probabilities for each observation based on regression coefficient estimates obtained without that observation; these values were 0.912, 0.909, 0.896, 0.867, and 0.833, respectively. We can see that penalizing based on predictor costs (e.g. with $b=0.15$) can lead to great cost reduction with very little loss in performance compared to the benefit-only model. Perhaps surprisingly, the model with cost equal to only \$$\$4.00$ per patient has a C-statistic of 0.849, which indicates that it might have some utility in a triage scenario. However, our method provides the ability for practitioners to see which predictors can supplement the model to improve diagnosis beyond easily-obtained predictors like age and sex and a subjective pain rating from the patient.
Figure 4:
@@ 3.5 Computation
Computations for the case study in Section 3.4 were completed using a $2\times\text{E}5-2683\text{v}42.1$GHz (Broadwell) CPU supercomputer from Advanced Research Computing at Virginia Tech. For a single value of $b$, it takes 16 minutes and 40 seconds to enumerate all models and calculate posterior model probabilities (benefit-only or cost-adjusted) for all $2^{13}$ candidate models.
@@ 4 Discussion
We have presented an approach to adjust cost-penalizing model priors for cost-adjusted Bayesian model selection. Our approach extends the well-established FND prior (Fouskakis et al., 2009a) by giving the practitioner the ability to adjust the level of cost penalization on candidate predictors and visualize the resulting cost-adjusted posterior inclusion probabilities. We proposed functions, according to a tuning parameter $b$, of the ratios of marginal predictor costs relative to a baseline cost. The resulting ECP prior that we study provides adjusted levels of cost penalization controlled by the practitioner via the tuning parameter. The properties that our cost ratio functions adhere to ensure that our adjusted priors can easily reproduce a benefit-only analysis and allow for unit conversion without changing model selection results, making it useful for costs measured in a variety of ways. We have shown, through simulation, that adjusting the cost penalty according to our proposed functions helps to maintain the cost penalization for larger sample sizes. Our inclusion path approach, which plots the change in cost-adjusted posterior inclusion probabilities together across a range of adjusted cost penalties, provides a visual tool to learn the relative importance of predictors when accounting for their costs and to make decisions about individual predictors to meet an overall budget. Our method can be applied to any binary outcome (e.g. diagnosis) where medical practitioners need to make a decision or prediction based on predictors with quantifiable costs.
@@
This work extends the utility of the FND prior by adjusting the penalization on costly predictors in model selection. For example, suppose that the model selected using the FND prior has a total predictor cost that exceeds the practitioner’s or hospital’s designated budget. Then our proposed inclusion path can easily be used to see which predictor(s) have probabilities that fall below the desired threshold as the practitioner slides $b$ towards higher values. Similarly, if the practitioner seeks a higher-performing model that still penalizes candidate predictor costs to some extent, they can slide $b$ down, closer to 00, to learn which additional predictor(s) will improve model fit without causing an undue increase in the cost per observation. We applied our approach to a data set of $297$ heart disease patients and found that decreasing the cost penalization from the FND prior helps to identify predictors that can improve the diagnosis of heart disease while still appropriately penalizing the most costly predictors. The ECP provides a useful option for adjusting the magnitude of cost penalization. The matter of cost can implicitly raise ethical concerns, especially in the case of medical applications that directly impact the health and well-being of patients. Medical practitioners would be well-positioned to apply our methods, as they have the knowledge and experience to understand how collecting different predictors may impact patients’ health, comfort, and finances and can advise on setting the cost ratio tuning parameter with these considerations in mind. We envision practitioners using a range of $g(\cdot)$ and $b$ values to study different cost penalizing settings and choose a model that best suits their budget and concerns.
@@
Avenues for future research include developing cost-penalized model selection for a sequential decision-making framework, for example, in medical diagnoses where practitioners may order additional tests depending on a patient’s initial results. Investigation may also be done to find and recommend an upper bound for $b$ given a particular data set. Our proposed method may also be adapted and applied more broadly to non-binary response data, for example to other generalized linear models. Another possible extension could include adjusting the model space and/or prior to accommodate cost structures for grouped or discounted predictors.

### v3

@@ 1 Introduction
When applying the FND prior to other data sets, we have found that the FND prior can lead to selection of less costly models when predictor costs differ, but the cost penalization is not appropriate or sufficient for all sample sizes. In fact, at large sample sizes, the cost penalization imposed by the FND prior greatly diminishes, often causing the resulting cost-penalized model selection to closely resemble that of a standard benefit-only analysis. To establish this phenomenon, we calculated the Kullback-Leibler (KL) divergence, which measures the difference between two probability distributions, between the sets of posterior model probabilities for all candidate models produced by cost-penalized analysis with the FND prior and the benefit-only selection approach as sample size increases for several simulated data sets. Since we consider the case of linear logistic regression, for which there is no closed form integrated likelihood, we examine the KL divergence comparing posterior model probabilities from the two approaches, where all posterior model probabilities are obtained after applying a Laplace approximation to the integrated likelihoods. We generated $10$ data sets of size $n=150$ with $p=9$ total candidate predictors with different costs as detailed in Section 3.1 and calculated the KL divergence between the two sets of posterior model probabilities produced by the benefit-only and FND selection approaches for each data set. Then, to imitate collection of additional data, we recursively added $300$ observations to the initial $10$ data sets, obtaining model selection results and calculating KL divergence between the posterior model probabilities produced by the two selection approaches for each new (larger) data set. Figure 2 plots the KL divergence between the posterior model probabilities produced by the two approaches as sample size increases for the $10$ collections of data sets increasing in size.
@@ 2.2 Cost-penalizing model selection
Following Fouskakis et al. (2009a), we consider the linear logistic regression setting in this paper, later applying it to the binary diagnosis of heart disease. For convenience, we use the notation of Fouskakis et al. (2009a) for indicator functions and predictor costs. We use a linear logistic regression model where $Y_{i}\in\{0,1\}$ and $X_{ij}$ denotes the $j^{th}$ predictor for observation $i$, where $i=1,\ldots,n$ and $j=0,\ldots,p$. Let $\gamma_{j}$ be an indicator that is equal to $1$ if predictor $X_{j}$ is included in the model and 00 if it is not. Then $\boldsymbol{\gamma}$ is a length $p$ vector of 00’s and $1$’s indicating whether each of the $p$ predictors are in the model or not. Note that the intercept is included in every model so $X_{i0}=1$ and $\gamma_{0}=1$ for all models. Then the Bayesian modeling framework is
@@
Equation (6) uses the marginal cost $\{c_{j},j=1,\ldots p\}$ of each candidate predictor, which is the cost of collecting that predictor for a single observation. When adhering to an overall budget, the practitioner may also want to consider the total cost per observation or the total cost of the model. The total cost per observation is equal to the sum of the marginal cost of all predictors in the model, i.e. $\sum_{j=1}^{p}c_{j}\gamma_{j}$, and the total cost of the model for the given observations is $n\times\sum_{j=1}^{p}c_{j}\gamma_{j}$.
@@ 2.3 Adjusted cost-penalizing functions
We propose functions $g(c_{j}/c_{0},b)$ to adjust the cost ratio according to tuning parameter $b$ that satisfy the following properties:
@@
    $b=0$ implies $g(c_{j}/c_{0},b)=1$ for all $j=1,\ldots,p$, reducing the prior to the uniform model prior (7) and resulting in a benefit-only model selection.
@@
    $b=1$ makes $g(c_{j}/c_{0},b)$ equal to $c_{j}/c_{0}$ and reproduces the FND prior as seen in Equation (6).
@@
    When $b>1$, $g(c_{j}/c_{0},b)$ penalizes predictors with costs $c_{j}>c_{0}$ more highly than the FND prior in (6). When $10<b<$ 1, $g(c_{j}/c_{0},b)$ penalizes predictors with costs $c_{j}>c_{0}$ less than the FND prior but more than a benefit-only analysis. Higher values of $b$ increase $g(c_{j}/c_{0},b)$ and the resulting cost penalization, leading to lower cost-adjusted prior inclusion probabilities for predictors with costs above the baseline.
@@
    When $c_{j}=c_{0}$, the $j^{th}$ candidate predictor cost is the same as the baseline cost. Changing $b$ does not introduce/increase penalization on any of the predictors with baseline cost.
@@ 2.4 Inclusion paths
Using the ECP and varying $b$, we create an inclusion path for each candidate predictor. To do this, we compute and plot cost-adjusted posterior inclusion probabilities for each of the candidate predictors across multiple values of $b$. See the Supplementary Material for detailed steps for constructing the inclusion path. The inclusion path visualizes a probabilistic way to learn the order in which the candidate predictors would be chosen or discarded, e.g. according to some threshold, if a different magnitude of cost penalization is used. The costs $\{c_{j},j=1,\ldots p\}$ remain unchanged; $g(c_{j}/c_{0},b)$ changes how heavily candidate predictors with larger costs are penalized a priori. Our inclusion path technique is inspired by the interpretable path diagrams like LASSO and ridge (Tibshirani, 1996; Hoerl and Kennard, 1970). From the plot of inclusion paths, the practitioner can all at once study candidate predictors’ posterior inclusion probabilities for the benefit-only analysis and cost-adjusted posterior inclusion probabilities for the FND prior and a wide range of the ECP with differing $b$ values. This is because we formulated the ECP to satisfy properties (a)-(d) so that the established uniform priors and the FND prior can be easily recreated from the ECP.
@@ 3.3 Inclusion paths using adjusted cost penalization
From Figure 3(a), the posterior inclusion probabilities for $8$ of the $9$ predictors decrease as the cost penalization is increased via larger values of $b$. The predictor with the larger effect size and baseline cost maintains a posterior inclusion probability (benefit-only or cost-adjusted) at $1$ across all values of $b$, as it is not penalized a priori by the FND or ECP priors, according to property (d) from Section 2.3. The posterior inclusion probabilities for the cheap and expensive null predictors are $0.09$ and $0.25$, respectively, in the benefit-only analysis corresponding to $b=0$. The values of $10<b<$ 1 introduce cost penalization that is less severe than that from the FND prior but higher than the (nonexistent) cost penalization from the benefit-only analysis. For example, the cost-adjusted posterior inclusion probabilities for the cheap and expensive null predictors drop to $0.04$ and $0.06$ when $b=0.2$ is used with the ECP. Then the null predictor with expensive cost has 00 cost-adjusted posterior inclusion probability for $b\geq 0.7$. Thus, even slightly penalizing these null predictors for having costs above the baseline helps to move their inclusion probabilities to 00. For $b=1$ (the FND approach), the null predictor with cheap cost has cost-adjusted posterior inclusion probability $0.0002$, which decreases to 00 for $b\geq 1.1$. The posterior inclusion probability of the null predictor with baseline cost is $0.12$ in the benefit-only analysis and the cost-adjusted posterior inclusion probability equals $0.05$ for $b\geq 3.5$, as this predictor has a null effect but is not penalized by any variation of the ECP. When determining which predictors are selected with each variation of the ECP, we can consider comparing the cost-adjusted posterior inclusion probabilities for the remaining non-null predictors to a threshold, e.g. $0.5$ as in the median probability model (Barbieri and Berger, 2004). The inclusion path allows us to see which predictors’ cost-adjusted inclusion probabilities meet the desired threshold for the different versions of cost-penalized selection at one time. For example, in the benefit-only analysis, which corresponds to the ECP with $b=0$, every one of the $6$ non-null predictors have posterior inclusion probabilities beyond the $0.5$ threshold under consideration. Using the FND prior, which corresponds to the ECP with $b=1$, the cost-adjusted posterior inclusion probabilities for all non-null predictors except for the expensive one with smaller effect size exceed the $0.5$ threshold. The FND approach is model selection consistent. Thus, for our simulated examples where $p=9$ and the full model space can be enumerated, the FND approach selects the true generating model with non-null predictors as the sample size increases. Figure 3 indicates that the cost-adjusted posterior inclusion probabilities fall below 0.5 at b = 1, 1.1, 1.5, and 1.8 for these predictors in the following order: expensive with smaller effect size, expensive with larger effect size, cheap with smaller effect size, and cheap with larger effect size.
@@
Figure 3(b) plots the cost of the selected model for each value of $b$ used to create the inclusion path in Figure 3(a). For this data set, the MAP model and median probability model are the same for all values of $b$. We can see that the cost of the selected model drops each time a predictor is no longer in the model, e.g. inclusion falls below 0.5. Figure 3(c) plots the accuracy in the form of the C-statistic of the selected model for each value of $b$. Similar to the cost in Figure 3(b), the C-statistic drops each time a predictor leaves the selected model as a result of a higher cost penalty. A practitioner, patient, or other stakeholder may use Figures 3(b) and 3(c) in conjunction with the inclusion path so they may keep any necessary budget, or desired efficiency or accuracy in mind when selecting cost-effective predictors.
@@ 3.4 Case study: selecting cost-effective predictors to model diagnosis of heart disease
We performed Bayesian model selection using the heart disease data, applying a benefit-only approach and cost-penalized model selection as described in Section 2 with the FND and ECP priors assigned to the model space. Benefit-only model selection with uniform priors (7) on the model space selects the median probability model with the following predictors: sex, chest pain, resting BP, ST depression, peak ST segment, major vessels, and defect type with corresponding posterior model probability $0.072$, up from a prior model probability equal to 0.000122 and with a total cost of \$$\$381.40$ per patient. Model selection using the FND prior on the model space selects the median probability model with the four baseline predictors (sex, age, chest pain, and resting BP), with corresponding cost-adjusted posterior model probability equal to 0.5. This model that contains only baseline predictors has prior model probability equal to $0.063$ using the FND prior, and a cost equal to \$$\$4.00$ per patient. While cost-penalized selection with the FND prior leads to a less costly model that is only $2.3\%$ of the cost of the benefit-only model, this model performs worse than more costly models in terms of classification. Figure 4 displays ROC curves and corresponding costs for models selected using different values of $b$ with ECP; the values of $b$ listed correspond to the median probability model. Detection of a health condition such as heart disease is critical for doctors to be able to provide effective treatment and advice to affected patients. The cost \$$\$4.00$ model might not have acceptable false positive rates for the hospital to trust the diagnosis, and thus it is desirable to be able to adjust the cost penalization, particularly between $10<b<$ 1, for these data. Our method is key for striking a balance between the benefit-only approach and the FND approach, by furnishing a spectrum of cost penalization choices. The C-statistics for the models considered in Figure 4 are 0.934, 0.931, 0.917, 0.887, and 0.849 for the median probability models from the ECP with $b=$ 0, 0.15, 0.25, 0.35, and 1, respectively. To assess out-of-sample predictive accuracy, we also calculated an average C-statistic based on 10-fold cross-validation for each model by obtaining predicted probabilities for each observation based on regression coefficient estimates obtained without each of the 10 folds; these values were 0.914, 0.913, 0.902, 0.86, and 0.833, respectively. See the Supplementary Material for more details. We can see that penalizing based on predictor costs (e.g. with $b=0.15$) can lead to great cost reduction with very little loss in performance compared to the benefit-only model. Perhaps surprisingly, the model with cost equal to only \$$\$4.00$ per patient has a C-statistic of 0.849, which indicates that it might have some utility in a triage scenario. However, our method provides the ability for practitioners to see which predictors can supplement the model to improve diagnosis beyond easily-obtained predictors like age and sex and a subjective pain rating from the patient.
@@ 3.5 Computation
Computations for the case study in Section 3.4 were completed using a $2\times\text{E}5-2683\text{v}42.1$GHz (Broadwell) CPU supercomputer from Advanced Research Computing at Virginia Tech. For a single value of $b$, it takes 16 minutes and 40 seconds to enumerate all models and calculate posterior model probabilities (benefit-only or cost-adjusted) for all $2^{13}$ candidate models. The examples studied here contain relatively small numbers of candidate predictors, so the entire model space of size $2^{p}$ can be enumerated. In the case of large $p$, our method could be adapted for use with a stochastic search algorithm to feasibly explore the model space, such as genetic algorithms (Wu et al., 2020) or reversible jump (Fouskakis et al., 2009a, c).
@@ 4 Discussion
We have presented an approach to adjust cost-penalizing model priors for cost-adjusted Bayesian model selection. Our approach extends the well-established FND prior (Fouskakis et al., 2009a) by giving the practitioner the ability to adjust the level of cost penalization on candidate predictors and visualize the resulting cost-adjusted posterior inclusion probabilities. We proposed functions, according to a tuning parameter $b$, of the ratios of marginal predictor costs relative to a baseline cost. The resulting ECP prior that we study provides adjusted levels of cost penalization controlled by the practitioner via the tuning parameter. The properties that our cost ratio functions adhere to ensure that our adjusted priors can easily reproduce a benefit-only analysis and allow for unit conversion without changing model selection results, making it useful for costs measured in a variety of ways. We have shown, through simulation, that adjusting the cost penalty according to our proposed functions helps to maintain the cost penalization for larger sample sizes. Our inclusion path approach, which plots the change in cost-adjusted posterior inclusion probabilities together across a range of adjusted cost penalties, provides a visual tool to learn the relative importance of predictors when accounting for their costs and to make decisions about individual predictors to meet an overall budget. Our method can be applied to any binary outcome (e.g. diagnosis) where medical practitioners need to make a decision or prediction based on predictors with quantifiable costs. Our proposed method does not have all of the same theoretical properties derived in Fouskakis et al. (2009a); for example, our proposed model prior is not model selection consistent. Rather, the purpose of our method is to enable the user to explore different cost penalty scenarios.
@@
This work extends the utility of the FND prior by adjusting the penalization on costly predictors in model selection. For example, suppose that the model selected using the FND prior has a total predictor cost that exceeds the practitioner’s or hospital’s designated budget (Fouskakis et al., 2009c). Then our proposed inclusion path can easily be used to see which predictor(s) have probabilities that fall below the desired threshold as the practitioner slides $b$ towards higher values. Similarly, if the practitioner seeks a higher-performing model that still penalizes candidate predictor costs to some extent, they can slide $b$ down, closer to 00, to learn which additional predictor(s) will improve model fit without causing an undue increase in the cost per observation. We applied our approach to a data set of $297$ heart disease patients and found that decreasing the cost penalization from the FND prior helps to identify predictors that can improve the diagnosis of heart disease while still appropriately penalizing the most costly predictors. The ECP provides a useful option for adjusting the magnitude of cost penalization. It is possible that, due to the presence of collinearity, large $p$, or other data characteristics, Bayesian model selection when the sample size is large may select different benefit-only and cost-penalized models with no adjustment to the cost penalty. For example, Fouskakis et al. (2009a) studied pneumonia-related deaths in $n=2,532$ patients with 83 candidate predictors, and cost-adjusted BIC selected a model which resulted in a 66.7% reduction in cost compared to the benefit-only model. Our method is intended to provide an extension for applications and data sets where the sample size impacts cost-penalized selection decisions, and to allow practitioners to consider how influential the prior should be with respect to their aim.
The matter of cost can implicitly raise ethical concerns, especially in the case of medical applications that directly impact the health and well-being of patients. Medical practitioners would be well-positioned to apply our methods, as they have the knowledge and experience to understand how collecting different predictors may impact patients’ health, comfort, and finances and can advise on setting the cost ratio tuning parameter with these considerations in mind. We envision practitioners using a range of $g(\cdot)$ and $b$ values to study different cost penalizing settings and choose a model that best suits their budget and concerns. We recommend that practitioners and researchers apply our cost-penalized selection methods in an ethical and transparent manner and acknowledge any implications of analyzing a particular type of cost.
@@
Avenues for future research include developing cost-penalized model selection for a sequential decision-making framework, for example, in medical diagnoses where practitioners may order additional tests depending on a patient’s initial results. Investigation may also be done to find and recommend an upper bound for $b$ given a particular data set. Our proposed method may also be adapted and applied more broadly to non-binary response data, for example to other generalized linear models. Another possible extension could include adjusting the model space and/or prior to accommodate cost structures for grouped or discounted predictors. Future development of an MCMC algorithm would allow for study of additional metrics of predictive ability based on the posterior distribution. Another opportunity for exploration may include changes to the prior and selection for categorical predictors such as those in Table 4. García-Donato and Paulo (2022) find that selecting either all or none of the levels of a categorical predictor can have disadvantages such as understating the effect of a categorical predictor with many levels if only a few of those levels are relevant. García-Donato and Paulo (2022) propose to allow for selection of only active levels of a categorical predictor and propose a model prior which maintains $1/2$ prior inclusion for categorical variables regardless of the number of levels. Further investigation is needed to understand how to apply this approach while incorporating a cost penalty on predictors through the model prior.